# The non-existence of some Galois representations of moderate degree in small characteristic

### v1

@@ 6. Non-existence of ${\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$-valued representations (Thm D, Cor E)
    $G\subseteq{\rm SO}^{+}(\mathbb{F}_{2^{r}})$.
    It is known (cf. [BHRD, Section 7.2.2]) that ${\rm SO}^{+}(\mathbb{F}_{2^{r}})$ has an index $2$ subgroup isomorphic to
    ${\rm PGL}_{2}(\mathbb{F}_{2^{r}})\times{\rm PGL}_{2}(\mathbb{F}_{2^{r}})$ $={\rm PSL}_{2}(\mathbb{F}_{2^{r}})\times{\rm PSL}_{2}(\mathbb{F}_{2^{r}})$
    $={\rm SL}_{2}(\mathbb{F}_{2^{r}})\times{\rm SL}_{2}(\mathbb{F}_{2^{r}}).$
    Thus, there exist an extension $L/\mathbb{Q}$ of degree at most $2$, and two Galois representations $\tau_{i}:G_{L}\longrightarrow{\rm SL}_{2}(\mathbb{F}_{2^{r}})\subset{\rm GL}_{2}(\mathbb{F}_{2^{r}})$ such that $\overline{\rho}|_{G_{L}}\simeq\tau_{1}\otimes\tau_{2}$. Since $\overline{\rho}$ is unramified outside $2$,
@@
    $\tau_{1}(g)={\rm diag}(a,a),\quad\tau_{2}(g)={\rm diag}(a^{-1},a^{-1})\quad\text{for some }a\in\mathbb{F}^{\times}_{2^{r}}.$
    But ${\rm Im}(\tau_{i})\subset{\rm SL}_{2}(\overline{{\mathbb{F}}}_{2^{r}})$, so we conclude that $\tau_{i}(I_{v})=\{1\}$ for $i=1,2$. Hence $\tau_{i}$ is unramified outside $2$. Applying [MT], $\tau_{i}$ has to be reducible. Since $G_{L}$ is normal in $G_{\mathbb{Q}}$, by Clifford’s theorem, $\overline{\rho}|_{G_{L}}$ is isomorphic to the direct sum of four one-dimensional representations of $G_{L}$ that are unramified outside $2$. However, all of them are the trivial representation of $G_{L}$ by Lemma 6.1. Thus, $|{\rm Im}(\overline{\rho})|\leqslant 2$ and $\overline{\rho}$ can not be irreducible over $\overline{{\mathbb{F}}}_{2}$. We conclude that this case cannot occur.
@@
    $G\subseteq{\rm SO}_{4}^{-}(\mathbb{F}_{2^{r}})$.
    It is known (cf. [BHRD, Section 7.2.2]) that ${\rm SO}^{-}(\mathbb{F}_{2^{r}})$ has an index $2$ subgroup isomorphic to
    ${\rm PGL}_{2}(\mathbb{F}_{2^{2r}})={\rm PSL}_{2}(\mathbb{F}_{2^{2r}})={\rm SL}_{2}(\mathbb{F}_{2^{2r}}).$
    Thus, there exist an extension $L/\mathbb{Q}$ of degree at most $2$ unramified outside $2$ and a continuous group homomorphism $\overline{\rho}|_{G_{L}}:G_{L}\longrightarrow{\rm GL}_{2}(\mathbb{F}_{2^{2r}})$. As in the previous case, $\overline{\rho}$ cannot be irreducible over $\overline{{\mathbb{F}}}_{2}$. We conclude that this case cannot occur.
@@
    $G\subseteq{\rm Sz}(\mathbb{F}_{2^{r}})$.
    In the above classification of the maximal subgroups of ${\rm Sz}(\mathbb{F}_{2^{r}})$, cases (Sz-1)–(Sz–3) cannot occur since $\overline{\rho}$ is irreducible over $\overline{{\mathbb{F}}}_{2}$.
@@
    The only possibility is therefore (Sz-5): $G={\rm Sz}(\mathbb{F}_{2^{r}})$ for some odd integer $r$. But $G\subseteq{\rm Sz}(\mathbb{F}_{2})=C_{5}\rtimes C_{4}$ (the Frobenius group of order $20$) can be seen to be impossible by the same argument as above, since it contains the dihedral group $D_{10}$ of order $10$.
The last claim in the statement of Theorem D follows from the fact that $|{\rm Sz}(\mathbb{F}_{2^{3}})|=29120$ and $|{\rm Sp}_{4}(\mathbb{F}_{2^{2}})|=979000$. ∎
@@
Assume GRH. There exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$.
@@
$\mathbb{Q}_{\overline{\rho}}:\mathbb{Q}]=|{\rm Im}(\overline{\rho})|<4800.$

### v2

@@ 6. Non-existence of $\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$-valued representations (Thm D, Cor E)
    $G\subseteq\mathrm{SO}^{+}(\mathbb{F}_{2^{r}})$.
    It is known (cf. [BHRD, Section 7.2.2]) that $\mathrm{SO}^{+}(\mathbb{F}_{2^{r}})$ has an index $2$ subgroup isomorphic to
    $\mathrm{PGL}_{2}(\mathbb{F}_{2^{r}})\times\mathrm{PGL}_{2}(\mathbb{F}_{2^{r}})$ $=\mathrm{PSL}_{2}(\mathbb{F}_{2^{r}})\times\mathrm{PSL}_{2}(\mathbb{F}_{2^{r}})$
    $=\mathrm{SL}_{2}(\mathbb{F}_{2^{r}})\times\mathrm{SL}_{2}(\mathbb{F}_{2^{r}}).$
    Thus, there exist an extension $L/\mathbb{Q}$ of degree at most $2$, and two Galois representations $\tau_{i}:G_{L}\longrightarrow\mathrm{SL}_{2}(\mathbb{F}_{2^{r}})\subset\mathrm{GL}_{2}(\mathbb{F}_{2^{r}})$ such that $\overline{\rho}|_{G_{L}}\simeq\tau_{1}\otimes\tau_{2}$. Since $\overline{\rho}$ is unramified outside $2$,
@@
    $\tau_{1}(g)=\mathrm{diag}(a,a),\quad\tau_{2}(g)=\mathrm{diag}(a^{-1},a^{-1})\quad\text{for some }a\in\mathbb{F}^{\times}_{2^{r}}.$
    But $\mathrm{Im}(\tau_{i})\subset\mathrm{SL}_{2}(\overline{\mathbb{F}}_{2^{r}})$, so we conclude that $\tau_{i}(I_{v})=\{1\}$ for $i=1,2$. Hence $\tau_{i}$ is unramified outside $2$. Applying [MT], $\tau_{i}$ has to be reducible. Since $G_{L}$ is normal in $G_{\mathbb{Q}}$, by Clifford’s theorem, $\overline{\rho}|_{G_{L}}$ is isomorphic to the direct sum of four one-dimensional representations of $G_{L}$ that are unramified outside $2$. However, all of them are the trivial representation of $G_{L}$ by Lemma 6.1. Thus, $|\mathrm{Im}(\overline{\rho})|\leqslant 2$ and $\overline{\rho}$ can not be irreducible over $\overline{\mathbb{F}}_{2}$. We conclude that this case cannot occur.
@@
    $G\subseteq\mathrm{SO}_{4}^{-}(\mathbb{F}_{2^{r}})$.
    It is known (cf. [BHRD, Section 7.2.2]) that $\mathrm{SO}^{-}(\mathbb{F}_{2^{r}})$ has an index $2$ subgroup isomorphic to
    $\mathrm{PGL}_{2}(\mathbb{F}_{2^{2r}})=\mathrm{PSL}_{2}(\mathbb{F}_{2^{2r}})=\mathrm{SL}_{2}(\mathbb{F}_{2^{2r}}).$
    Thus, there exist an extension $L/\mathbb{Q}$ of degree at most $2$ unramified outside $2$ and a continuous group homomorphism $\overline{\rho}|_{G_{L}}:G_{L}\longrightarrow\mathrm{GL}_{2}(\mathbb{F}_{2^{2r}})$. As in the previous case, $\overline{\rho}$ cannot be irreducible over $\overline{\mathbb{F}}_{2}$. We conclude that this case cannot occur.
@@
    $G\subseteq\mathrm{Sz}(\mathbb{F}_{2^{r}})$.
    In the above classification of the maximal subgroups of $\mathrm{Sz}(\mathbb{F}_{2^{r}})$, cases (Sz-1)–(Sz–3) cannot occur since $\overline{\rho}$ is irreducible over $\overline{\mathbb{F}}_{2}$.
@@
    The only possibility is therefore (Sz-5): $G=\mathrm{Sz}(\mathbb{F}_{2^{r}})$ for some odd integer $r$. But $G\subseteq\mathrm{Sz}(\mathbb{F}_{2})=C_{5}\rtimes C_{4}$ (the Frobenius group of order $20$) can be seen to be impossible by the same argument as above, since it contains the dihedral group $D_{10}$ of order $10$.
The last claim of Theorem D follows from the fact that $|\mathrm{Sz}(\mathbb{F}_{2^{3}})|=29120$ and $|\mathrm{Sp}_{4}(\mathbb{F}_{2^{2}})|=979000$. ∎
@@
Assume GRH. There exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$.
@@
$\mathbb{Q}_{\overline{\rho}}:\mathbb{Q}]=|\mathrm{Im}(\overline{\rho})|<4800.$